OpenStudy (anonymous):

solve sec x = -2, finding All solutions. express the solutions in either radians or degrees.

OpenStudy (anonymous):

\[\frac{1}{\cos(x)}=-2\iff \cos(x)=-\frac{1}{2}\] and work from there

OpenStudy (anonymous):

i.e. your job is to find all numbers \(x\) with \(\cos(x)=-\frac{1}{2}\)

OpenStudy (anonymous):

there are two places on the unit circle where the first coordinate is \(-\frac{1}{2}\) so first you need to find them, and then write a general formula that will give all of them second part is easy, take the answer to the first part and write \(+2k\pi\)